solve by the substitution method 4x-2y=-30 8x+81=y
Answer by stanbon(75887)

solve by the substitution method 4x-2y=-30 8x+81=y
------------------
Substitute for "y" in the 1st equation:
---
4x- 2(8x+81) = -30
4x-16x - 162 = -30
-12x = 132
x= -11
-----
Substitute into y = 8x+81 to solve for "y":
y = 8(-11)+81
y = -88+81
y = -7
==============
cheers,
Stan H.
==============